# Dissecting the Hubble tension: Insights from a diverse set of Sound Horizon-free measurements

### v1

@@ Abstract
The Hubble tension is commonly framed as a discrepancy between local, late-time measurements favoring km s-1 Mpc-1 and early-time, Sound-Horizon-based measurements favoring km s-1 Mpc-1. We challenge this viewpoint by analyzing 83 Sound-Horizon-independent measurements, categorized into four classes: Distance Ladder measurements using local calibrators; Local One-Step CDM measurements assuming the standard expansion history; Pure Local One-Step measurements independent of shape; and CMB Sound Horizon free measurements using CMB data without the Sound Horizon scale. Our analysis reveals that the 29 Distance Ladder measurements yield km s-1 Mpc-1 (), while the 54 One-Step measurements collectively yield km s-1 Mpc-1 (), a tension exceeding the Planck–SH0ES discrepancy. This tension remains significant at after accounting for correlations. Among One-Step categories, Local One-Step CDM measurements favor the lowest value ( km s-1 Mpc-1, ), Pure Local One-Step yield an intermediate value ( km s-1 Mpc-1, ), and CMB Sound Horizon Free measurements give km s-1 Mpc-1 (). We conclude that the Hubble tension is better characterized as a discrepancy between the Distance Ladder and all other methodologies, rather than an early-vs-late-time split. We also identify a internal tension among One-Step measurements: analyses assuming CDM systematically recover lower values by approximately 3.2 km s-1 Mpc-1 compared to model-independent methods. This suggests either unrecognized systematics/new physics in the Distance Ladder or deviations from CDM in the late-time Universe.
@@ I Introduction
The standard view frames this tension as a dichotomy between “early” and “late” times. The Planck CMB analysis, assuming CDM, yields [Planck2018], an inference dependent on the Sound Horizon scale at the drag epoch (). The local Distance Ladder, anchored by geometric measurements and extended via Cepheids to SNe Ia, favors [Riess2022], reinforced by JWST observations [Riess2025] and extended network analyses [Casertano2025]. If both are accurate, CDM is incomplete, potentially necessitating new physics [DiValentino2021review, Knox2020, Abdalla:2022yfr, Perivolaropoulos:2021jda, Alestas2020]. Recent DESI hints of dynamical dark energy evolving across the phantom divide [Efstratiou2025, Ozulker:2025ehg, Scherer:2025esj] further complicate this picture, with phantom crossings previously proposed to address both the Hubble tension and the SNe Ia absolute magnitude discrepancy [Alestas2021].
@@
However, this “Early vs. Late” characterization may be oversimplified. It assumes all local measurements favor high values, an assumption challenged by “One-Step” measurements independent of the Cepheid/SN Ia calibration chain. Our recent analysis [Perivolaropoulos2024] demonstrated a stark bifurcation: Distance Ladder measurements clustered around , while independent One-Step methods statistically favored lower values closer to Planck. This suggests the discrepancy may lie between the distance ladder methodology and other probes, rather than between early and late Universe physics [Huang:2024].Indeed, the tension has been shown to have a multidimensional character, with degeneracies between , , and other parameters playing a crucial role [Pedrotti:2024kpn].
@@
In this work, we present an up-to-date compilation of 83 measurements published through late-2025, all independent of the CMB Sound Horizon scale. We introduce a four-category classification based on methodology and model dependence:
1.
Category 1 – Distance Ladder (29 measurements): Traditional three-rung approach using geometric anchors to calibrate standard candles (Cepheids, TRGB, JAGB, Miras) that standardize SNe Ia in the Hubble flow.
2.
Category 2 – Local One-Step CDM (33 measurements): Local measurements bypassing the Distance Ladder but assuming CDM to infer (e.g., Gravitational Wave sirens, Strong Lensing time delays, FRB dispersion measures).
3.
Category 3 – Pure Local One-Step (15 measurements): Model-independent local measurements without assuming a specific shape (e.g., geometric Megamaser distances, Cosmic Chronometers, cosmographic lensing analyses).
4.
Category 4 – CMB Sound Horizon Free (6 measurements): CMB-based measurements decoupled from , using alternative scales such as the horizon at matter-radiation equality ().
@@
The classification follows a systematic decision process (Figure 1): measurements using multi-rung calibration chains are assigned to Category 1; those utilizing CMB data without to Category 4; remaining measurements are assigned to Category 2 or 3 based on whether they assume CDM. The same observational technique can yield measurements in different categories depending on analysis methodology—for example, Megamaser observations with Keplerian models belong to Category 3 [Kuo2013], while those adopting CDM priors belong to Category 2 [Barua2025].
@@
This classification isolates two distinct tensions: the external tension between the Distance Ladder and all other methods, and an internal tension within One-Step measurements examining whether CDM assumptions systematically shift inferred values.A comprehensive overview of observational tensions and potential resolutions involving both systematics and fundamental physics has been presented in the CosmoVerse White Paper [CosmoVerseNetwork:2025alb].
@@ II Data Compilation and Categorization
To dissect the Hubble tension beyond the traditional “Early vs. Late” dichotomy, we have compiled a comprehensive dataset of 83 independent measurements of the Hubble constant, updated through late-2025. This compilation builds upon the catalog presented in our previous work [Perivolaropoulos2024], significantly expanded to include recent results from James Webb Space Telescope (JWST) calibration campaigns, the third generation of gravitational wave catalogs (GWTC-3), and the latest releases from the Dark Energy Spectroscopic Instrument (DESI) [DESI:2025zgx].
@@ II.1 Selection Criteria
Independence of Estimates: Where multiple analyses utilize the exact same dataset (e.g., multiple re-analyses of the same supernova catalog with identical calibration), we select the most recent or representative result to avoid double-counting, unless the methodologies differ significantly in their systematic assumptions.
@@ II.2 Classification Methodology
Figure 1 illustrates the classification decision tree schematically. For each measurement, we first assess whether it employs a multi-rung calibration chain (leading to Category 1). If not, we determine whether it utilizes CMB data (leading to Category 4 if Sound-Horizon-free). For the remaining local One-Step measurements, we examine whether the analysis assumes a CDM expansion history (Category 2) or proceeds model-independently (Category 3).
@@
| Technique | Category 2 (CDM) | Category 3 (Model-Indep.) |
@@ II.3 Category 1: Distance Ladder Measurements
This category comprises 29 measurements that employ the traditional multi-rung calibration method. These measurements anchor the absolute magnitude of Type Ia supernovae (SNe Ia) in the Hubble flow using local standard candles calibrated by geometric distances. This is the “Late Universe” standard, relying on the fundamental assumption that the physics of calibrators (Cepheids, TRGB, JAGB, Miras) and SNe Ia are uniform across cosmic time and environment.
@@
Our compilation (Table 2) includes Cepheid-based measurements and recent JWST updates [Riess2025], TRGB-based measurements including CCHP results [Freedman2024], and alternative calibrators such as JAGB stars, Mira variables, Surface Brightness Fluctuations (SBF), and the Tully-Fisher relation. We also include the galaxy pairwise velocity method [Zhang2024], which, while methodologically distinct, relies on calibration from distance ladder measurements.
@@
| Huang [Huang2020] | 2019 | 73.3 | 4.0 | Mira |
| de Jaeger [deJaeger2020] | 2020 | 75.8 | 5.1 | Cepheids/TRGB + SNe II |
| Khetan [Khetan2021] | 2021 | 70.5 | 4.1 | SBF |
| Freedman [Freedman2021] | 2021 | 69.8 | 1.7 | TRGB (CCHP) |
| Kenworthy [Kenworthy2022] | 2022 | 73.1 | 2.5 | Two-rung ladder |
| Scolnic [Scolnic2024] | 2023 | 73.22 | 2.06 | TRGB (SH0ES calibrators) |
| Li [Li2024b] | 2024 | 74.7 | 3.1 | JAGB + SNe Ia |
| Lee [Lee2024] | 2024 | 67.8 | 2.7 | JAGB + SNe Ia |
| Jensen [Jensen2025] | 2025 | 73.8 | 2.4 | TRGB |
| Riess [Riess2025] | 2025 | 73.49 | 0.93 | Cepheids (JWST) |
| Bhardwaj [Bhardwaj2025] | 2025 | 73.06 | 2.6 | Mira |
| Newman [Newman2025] | 2025 | 75.3 | 2.9 | TRGB + SNe Ia |
We note that the Lee [Lee2024] JAGB measurement ( km s-1 Mpc-1) represents a notable outlier within this category, yielding a value consistent with Planck rather than the typical Distance Ladder result. This discrepancy with other JAGB analyses (e.g., Li [Li2024b]) appears to originate from differences in the JAGB zero-point calibration and photometric selection criteria, highlighting that even within a single indicator class, methodological choices can significantly impact the inferred .
II.4 Category 2: Local One-Step (CDM Assumption)
This category includes 33 measurements that bypass the Distance Ladder (“One-Step”) but infer by explicitly assuming the standard CDM expansion history:
@@ II.4 Category 2: Local One-Step (CDM Assumption)
Gravitational Wave Standard Sirens: Both bright sirens (with EM counterparts, e.g., GW170817 [Palmese2023]) and dark sirens (using galaxy catalogs for redshift inference [Abbott2021, Beirnaert2025]), where the luminosity distance is computed assuming CDM.
Strong Lensing Time Delays: Analyses where the cosmological model enters both the time-delay distance calculation and often the lens mass modeling [Shajib2023, Kelly2023, Birrer2025].
Fast Radio Bursts: Measurements using the dispersion measure–redshift relation interpreted under CDM assumptions for the intergalactic medium [Wu2022, PiratovMoreno2025].
@@
| Wu [Wu2022] | 2021 | 68.8 | 4.7 | FRBs |
| Shajib [Shajib2023] | 2023 | 77.1 | 7.2 | Lensing (TDCOSMO) |
| Napier [Napier2023] | 2023 | 74.1 | 8.0 | Cluster Lenses |
| Kelly [Kelly2023] | 2023 | 66.6 | 3.7 | Lensed SN Refsdal |
| Liu [liu2023h] | 2023 | 59.1 | 3.6 | Cluster Lensed Quasar |
| Martinez [Martinez2023] | 2023 | 74.0 | 11.2 | Time-Delay Cosmography |
| Pascale [Pascale2024] | 2024 | 71.8 | 8.7 | Lensed SN H0pe (JWST) |
| Yang [Yang2024] | 2024 | 74.0 | 7.4 | FRBs w/ Scattering |
| Piratova [PiratovMoreno2025] | 2025 | 65.1 | 7.4 | FRBs (MLE) |
| Beirnaert [Beirnaert2025] | 2025 | 79.0 | 9.0 | GW Dark Sirens |
| Birrer [Birrer2025] | 2025 | 71.6 | 3.6 | TDCOSMO Strong Lensing |
| Pierel [Pierel2025] | 2025 | 66.9 | 9.8 | SN Encore Time Delays |
II.5 Category 3: Pure Local One-Step (Model Independent)
This category represents the cleanest test of the Distance Ladder, as it comprises 15 measurements that determine geometrically or via direct measurement of the expansion rate, without assuming a specific functional form for . These measurements are either:
Purely geometric: Relying on Keplerian dynamics or light-travel-time arguments that are independent of the global cosmological model (e.g., Megamaser disk modeling [Kuo2013, Gao2015]).
Direct measurements: Cosmic chronometers measure directly from the differential aging of massive, passively evolving galaxies, providing model-independent determinations of the expansion rate at various redshifts [Moresco2023]. To extract , two main approaches have been employed: (i) Gaussian Process reconstruction of combined with a cosmographic expansion of the luminosity distance in the Hubble flow (), (ii) using cosmic chronometers to calibrate the H ii galaxies Hubble diagram, which in turn anchors Type Ia supernovae at low redshift via Hubble’s law [Zhang2022]. Both methods avoid assumptions about the underlying cosmological model beyond the FLRW metric.
Cosmographic approaches: Analyses that expand or in Taylor series around , avoiding assumptions about the energy content of the Universe [Du2023].
Expanding photosphere methods: Geometric distance determinations from supernovae using photospheric velocity and angular size evolution [Sneppen2023, Vogl2024].

### v2

# Dissecting the Hubble tension: Insights from a diverse set of Sound-Horizon-Free measurements

@@ Abstract
The Hubble tension is commonly framed as a discrepancy between local, late-time measurements favoring km s-1 Mpc-1 and early-time, Sound-Horizon-based measurements favoring km s-1 Mpc-1. We challenge this viewpoint by analyzing 88 Sound Horizon Free measurements, categorized into four classes: Distance Ladder measurements using local calibrators; Local CDM measurements assuming the standard expansion history; Pure Local measurements independent of shape; and CMB Sound–Horizon–Free measurements using CMB data without the Sound Horizon scale. Our analysis reveals that the 30 Distance Ladder measurements yield km s-1 Mpc-1 (), while the 58 Distance Ladder-Independent/Sound Horizon Free measurements collectively yield km s-1 Mpc-1 (), a tension exceeding the Planck–SH0ES discrepancy. This tension remains significant at a minimum value of after accounting for correlations. Among categories, Local CDM measurements favor the lowest value ( km s-1 Mpc-1), Pure Local yield an intermediate value ( km s-1 Mpc-1), and CMB Sound Horizon Free measurements give km s-1 Mpc-1. We conclude that the Hubble tension is better characterized as a discrepancy between the Distance Ladder and all other methodologies, rather than an early-vs-late-time split. We also identify a internal tension among Distance Ladder Independent/Sound Horizon Free measurements (noting its sensitivity to categorization choices): analyses assuming CDM systematically recover lower values by 3.4 km s-1 Mpc-1 compared to cosmological-model-independent methods. This suggests either unrecognized systematics in the Distance Ladder or deviations from CDM or both.
@@ I Introduction
The standard view frames this tension as a dichotomy between “early” and “late” times. The Planck CMB analysis, assuming CDM, yields [141], an inference dependent on the Sound Horizon scale at the drag epoch (). The local Distance Ladder, anchored by geometric measurements and extended via Cepheids to SNe Ia, favors [150], reinforced by JWST observations [149] and extended network analyses [31]. If both are accurate, CDM is incomplete, potentially necessitating new physics [51, 91, 3, 131, 9, 179, 33, 135].
Various theoretical approaches have been proposed to address the Hubble tension, which can be broadly classified into three categories based on the cosmic epoch at which they introduce new physics [51, 91].
Early-time models attempt to resolve the tension by modifying conditions during the recombination epoch, thereby reducing the Sound Horizon scale that serves as the standard ruler for CMB-based inference. Prominent examples include Early Dark Energy (EDE) scenarios and modified gravity theories. However, these approaches face significant challenges: they typically fail to fully resolve the tension, often exacerbate the tension by enhancing matter clustering, and require considerable fine-tuning of model parameters [153, 168, 146, 169, 170, 82, 181, 73].
Late-time models propose deformations in the Hubble expansion history at low redshifts , often through dynamical dark energy with an evolving equation of state [143, 35, 158, 124, 162, 59, 49, 50, 97, 187, 61]. To understand the implications, it is essential to first outline the theoretical classification of Dark Energy dynamics. The evolution of dark energy is characterized by its equation of state parameter, , defined as the ratio of its pressure to its energy density ().
In the standard CDM model, this parameter corresponds to a Cosmological Constant and is strictly fixed at . However, dynamical dark energy models are generally categorized into two distinct regimes based on deviations from this value: Quintessence (), which is typically modeled by a scalar field rolling down a potential, and the Phantom regime (), where the fluid exhibits super-negative pressure, leading to an increasing energy density over time [120, 166]
A dynamical evolution in which the parameter traverses the boundary is referred to as Phantom Crossing. This transition presents significant theoretical challenges; within the framework of simple perfect fluids or single scalar field models, crossing this boundary typically triggers gradient instabilities or introduces ghost degrees of freedom. Consequently, viable phantom crossing scenarios usually necessitate frameworks beyond standard General Relativity, such as Modified Gravity theories or Effective Field Theory (EFT) approaches, to stabilize the transition [57, 118, 166, 160]. Recent analyses incorporating DESI Data Release 2 (DR2) [4] combined with CMB observations favor a Dynamical Dark Energy scenario characterized by a specific phantom crossing behavior. Contrary to simple quintessence models, the data indicate a transition from a phantom-like phase () in the past to a quintessence-like phase () at late times, crossing the phantom divide at approximately , with a present-day value of .
Crucially, regarding the Hubble tension, this result presents a paradox: due to the geometric anti-correlation between and , a present-day value of implies a lower inferred value for the Hubble constant. Consequently, rather than bridging the gap towards the Distance Ladder value ( km s-1 Mpc-1), this specific form of dynamical dark energy risks exacerbating the discrepancy with the high Distance Ladder value. [126, 58, 114, 30, 100, 38, 151, 166, 19, 85]. However, it aligns remarkably well with intermediate values: recent work by Zhang et al. using an interacting dark energy model finds from CMBDESI DR2 data (without SH0ES priors), a value consistent with intermediate estimates between Planck and SH0ES [194]. If the DESI evidence for dynamical dark energy is confirmed, the observed shift between CDM-dependent (Category 2) and model-independent (Category 3) measurements transforms from a statistical curiosity into a physical necessity. In this scenario, the Pure Local measurements (Category 3), yielding km s-1 Mpc-1, would emerge as the only methodologically sound Distance Ladder Independent/Sound Horizon Free determinations. Conversely, the lower values found in Category 2 would be interpreted as an artifact of imposing an invalid CDM expansion history on local data. [57, 120, 160, 10, 165, 71, 118, 152, 159, 2].
Ultralate-time models propose environmental or physical changes at very low redshifts that could affect the calibration of distance ladder indicators. The G-step model [112, 154, 128, 133], which posits a transition in the effective gravitational constant at , represents a well-studied example. Related approaches involving screened fifth forces have been explored by Desmond et al. [44, 45, 76], who showed that such modifications could systematically affect Cepheid and TRGB distance indicators differently. Alternative explanations within this category include local void scenarios, which posit that our local environment is underdense relative to the cosmic average, leading to an apparent acceleration of the local expansion rate [89, 109, 27, 32, 29, 111, 84, 15]. While these models offer mechanisms to reconcile distance ladder and other measurements without modifying early-universe physics, they require fine-tuning and lack complete theoretical justification for the proposed transitions.
@@
However, this “Early vs. Late” characterization may be oversimplified. It assumes all local measurements favor high values, an assumption challenged by Distance Ladder Independent/Sound Horizon Free measurements independent of the Cepheid/SN Ia calibration chain. Our recent analysis [134] demonstrated a stark bifurcation: Distance Ladder measurements clustered around , while independent methods statistically favored lower values closer to Planck. This suggests the discrepancy may lie between the distance ladder methodology and other probes, rather than between early and late Universe physics [80]. Indeed, the tension has been shown to have a multidimensional character, with degeneracies between , , and other parameters playing a crucial role [127].
@@
In this work, we present an up-to-date compilation of 88 H0 measurements published through late-2025, all independent of the CMB Sound Horizon scale. We introduce a four-category classification based on methodology and model dependence: (1) Distance Ladder (30 measurements), (2) Local CDM (33 measurements), (3) Pure Local (cosmological-model-independent, (16 measurements), and (4) CMB Sound Horizon Free (CMB–SHF) (9 measurements). The detailed definitions and classification criteria are presented in Section II.2.
@@
The classification follows a systematic decision process (Figure 1): measurements using multi-rung calibration chains are assigned to Category 1; those utilizing CMB data without to Category 4; remaining measurements are assigned to Category 2 or 3 based on whether they assume CDM. The same observational technique can yield measurements in different categories depending on analysis methodology—for example, Megamaser observations with Keplerian models belong to Category 3 [95], while those adopting CDM priors belong to Category 2 [17].
@@
This classification isolates two distinct tensions: the external tension between the Distance Ladder and all other methods, and an internal tension within these measurements examining whether CDM assumptions systematically shift inferred values. A comprehensive overview of observational tensions and potential resolutions involving both systematics and fundamental physics has been presented in the CosmoVerse White Paper [52].
@@ II Data Compilation and Categorization
To dissect the Hubble tension beyond the traditional “Early vs. Late” dichotomy, we have compiled a comprehensive set of 88 independent measurements of the Hubble constant, updated through late-2025. This compilation builds upon the catalog presented in our previous work [134], significantly expanded to include recent results from James Webb Space Telescope (JWST) calibration campaigns, the third generation of gravitational wave catalogs (GWTC-3), and the latest releases from the Dark Energy Spectroscopic Instrument (DESI) [4].
@@ II.1 Selection Criteria
Independence of Estimates: Where multiple analyses utilize the exact same sample (e.g., multiple re-analyses of the same supernova catalog with identical calibration), we select the most recent or representative result to avoid double-counting, unless the methodologies differ significantly in their systematic assumptions.
@@ II.2 Classification Methodology
Figure 1 illustrates the classification decision tree schematically. For each measurement, we first assess whether it employs a multi-rung calibration chain (leading to Category 1). If not, we determine whether it utilizes CMB data (leading to Category 4 if Sound-Horizon-Free). For the remaining local measurements, we examine whether the analysis assumes a CDM expansion history (Category 2) or proceeds cosmological-model-independently (Category 3).
@@
Throughout this work, we use the term “cosmological-model-independent” (Category 3) to denote measurements that do not assume a specific functional form for the expansion history —in particular, they do not assume the standard CDM expansion history. This should be understood as independence from cosmological model assumptions rather than complete model independence, as all measurements necessarily rely on certain physical assumptions (e.g., standard astrophysics, the FLRW metric, general relativity). Similarly, we define Distance Ladder-Independent/Sound Horizon Free (DLI–SHF) measurements (Categories 2, 3, and 4) as any determination that bypasses the traditional three-rung Distance Ladder calibration chain (geometric anchors Cepheids/TRGB/JAGB/Miras SNe Ia in the Hubble flow), regardless of whether it probes early or late cosmic times. Some DLI–SHF methods may involve their own calibration procedures (e.g., using cosmic chronometers to calibrate the HII galaxy Hubble diagram), but these do not rely on the standard candle–supernova chain that defines the Distance Ladder.
| Technique | Category 2 (CDM) | Category 3 (Cosmological-Model-Independent) |
@@ II.3 Category 1: Distance Ladder Measurements
This category comprises 30 measurements that employ the traditional multi-rung calibration method. These measurements anchor the absolute magnitude of Type Ia supernovae (SNe Ia) in the Hubble flow using local standard candles calibrated by geometric distances. This is the “Late Universe” standard, relying on the fundamental assumption that the physics of calibrators (Cepheids, TRGB, JAGB, Miras) and SNe Ia are uniform across cosmic time and environment.
@@
Our compilation (Table 2) includes Cepheid-based measurements and recent JWST updates [149], TRGB-based measurements including CCHP results [62], and alternative calibrators such as JAGB stars, Mira variables, Surface Brightness Fluctuations (SBF), and the Tully-Fisher relation. We also include the galaxy pairwise velocity method [193], which, while methodologically distinct, relies on calibration from distance ladder measurements [178, 56, 157].
@@
| Huang [78] | 2019 | 73.3 | 4.0 | Mira |
| de Jaeger [42] | 2020 | 75.8 | 5.1 | Cepheids/TRGB + SNe II |
| Khetan [90] | 2021 | 70.5 | 4.1 | SBF |
| Freedman [63] | 2021 | 69.8 | 1.7 | TRGB (CCHP) |
| Kenworthy [88] | 2022 | 73.1 | 2.5 | Two-rung ladder |
| Scolnic [161] | 2023 | 73.22 | 2.06 | TRGB (SH0ES calibrators) |
| Li [98] | 2024 | 74.7 | 3.1 | JAGB + SNe Ia |
| Lee [96] | 2024 | 67.8 | 2.7 | JAGB + SNe Ia |
| Lapuente [155] | 2025 | 72.61 | 1.69 | TRGB+JAGB+Ceph.+SNe Ia (twins) |
| Jensen [83] | 2025 | 73.8 | 2.4 | TRGB |
| Riess [149] | 2025 | 73.49 | 0.93 | Cepheids (JWST) |
| Bhardwaj [20] | 2025 | 73.06 | 2.6 | Mira |
| Newman [117] | 2025 | 75.3 | 2.9 | TRGB + SNe Ia |
We note that the Lee [96] JAGB measurement ( km s-1 Mpc-1) represents a notable outlier within this category, yielding a value consistent with Planck rather than the typical Distance Ladder result. This discrepancy with other JAGB analyses (e.g., Li, Lapuente [98, 155]) appears to originate from differences in the JAGB zero-point calibration and photometric selection criteria, highlighting that even within a single indicator class, methodological choices can significantly impact the inferred . A critical recent addition to this category is the analysis by Ruiz-Lapuente et al. [155], which utilizes the “SNe Ia twins” method calibrated via both TRGB and JAGB. Notably, their specific calibration using JAGB yields km s-1 Mpc-1, while their TRGB determination gives km s-1 Mpc-1. In our compilation (Table 2), we have adopted the value km s-1 Mpc-1, which represents the weighted average across the different methodological combinations presented in their work. This independent confirmation of high values using alternative calibrators is particularly significant as it directly challenges previous lower estimates ( km s-1 Mpc-1) historically associated with the CCHP program. It suggests that the previously reported discrepancies between Cepheids and other local calibrators of the Distance Ladder may have been overstated, implying that unrecognized systematics in the specific calibration or sample selection of earlier CCHP analyses—rather than the indicators themselves—may account for the lower values. Consequently, this result reinforces the internal consistency of Category 1, further isolating the Distance Ladder methodology from the rest of the cosmological probes.
II.4 Category 2: Local (CDM Assumption)
This category includes 33 measurements that bypass the Distance Ladder but infer by explicitly assuming the standard CDM expansion history:
@@ II.4 Category 2: Local (CDM Assumption)
| Wu [189] | 2021 | 68.8 | 4.7 | FRBs |
| Shajib [163] | 2023 | 77.1 | 7.2 | Lensing (TDCOSMO) |
| Napier [116] | 2023 | 74.1 | 8.0 | Cluster Lenses |
| Kelly [87] | 2023 | 66.6 | 3.7 | Lensed SN Refsdal |
| Liu [104] | 2023 | 59.1 | 3.6 | Cluster Lensed Quasar |
| Martinez [113] | 2023 | 74.0 | 11.2 | Time-Delay Cosmography |
| Pascale [125] | 2024 | 71.8 | 8.7 | Lensed SN H0pe (JWST) |
| Yang [190] | 2024 | 74.0 | 7.4 | FRBs w/ Scattering |
| Piratova [140] | 2025 | 65.1 | 7.4 | FRBs (MLE) |
| Beirnaert [18] | 2025 | 79.0 | 9.0 | GW Dark Sirens |
| Birrer [175] | 2025 | 71.6 | 3.6 | TDCOSMO Strong Lensing |
| Pierel [138, 174] | 2025 | 66.9 | 9.8 | SN Encore Time Delays |
II.5 Category 3: Pure Local (Cosmological-Model Independent)
This category represents the cleanest test of the Distance Ladder, as it comprises 16 measurements that determine geometrically or via direct measurement of the expansion rate, without assuming a specific functional form for . These measurements are either:
Direct measurements: Cosmic chronometers measure directly from the differential aging of massive, passively evolving galaxies, providing cosmological-model-independent determinations of the expansion rate at various redshifts [115]. To extract , two main approaches have been employed: (i) Gaussian Process reconstruction of combined with a cosmographic expansion of the luminosity distance in the Hubble flow (), (ii) using cosmic chronometers to calibrate the H ii galaxies Hubble diagram, which in turn anchors Type Ia supernovae at low redshift via Hubble’s law [192]. Both methods avoid assumptions about the underlying cosmological model beyond the FLRW metric.
Cosmographic approaches: Analyses that expand or in Taylor series around , avoiding assumptions about the energy content of the Universe [55].